### v2

# Partitioned Quantum Subspace Expansion

Abstract
We present an iterative generalisation of the quantum subspace expansion algorithm used with a Krylov basis. The iterative construction connects a sequence of subspaces via their lowest energy states. Diagonalising a Hamiltonian in a given Krylov subspace requires the same quantum resources in both the single step and sequential cases. We propose a variance-based criterion for determining a good iterative sequence and provide numerical evidence that these good sequences display improved numerical stability over a single step in the presence of finite sampling noise. Implementing the generalisation requires additional classical processing with a polynomial overhead in the subspace dimension. By exchanging quantum circuit depth for additional measurements the quantum subspace expansion algorithm appears to be an approach suited to near term or early error-corrected quantum hardware. Our work suggests that the numerical instability limiting the accuracy of this approach can be substantially alleviated beyond the current state of the art.
1 Introduction
Finding accurate ground state energies of complex quantum systems is of paramount importance for fundamental research and applied chemistry applications. However, the exponential scaling of the Hilbert space dimension with system size means that many problems of interest cannot be solved by classical computers. On the other hand, quantum computers can in principle address this problem by storing and processing complex quantum states with $2^{n}$ degrees of freedom using $n$ qubits.
Figure 1: Illustration of a Partitioned Quantum Subspace Expansion (PQSE) for $P=2$ partitions. |$\ket{\text{GS}}$ denotes the ground state of a Hamiltonian $H$. When diagonalising $H$ in the order $R$ Krylov subspace $\mathcal{K}_{R}(H,\ket{\phi_{0}})$, numerical instabilities can prohibit access to the ground state estimate |$\ket{\psi_{R}}$. PQSE can avoid this by carrying out two sequential steps: First we diagonalise $H$ in $\mathcal{K}_{r_{1}}(H,\ket{\phi_{0}})$ with $r_{1}<R$. We then use the lowest energy state |$\ket{\psi_{1}}$ of $H$ in $\mathcal{K}_{r_{1}}(H,\ket{\phi_{0}})$ to construct a second subspace $\mathcal{K}_{r_{2}}(H,\ket{\psi_{1}})$, where $r_{1}+r_{2}-1=R$, with lowest energy state |$\ket{\psi_{2}}$. For certain values of $\{r_{1},r_{2}\}$, the ground state estimate |$\ket{\psi_{2}}$ avoids the numerical instabilities which prohibit access to |$\ket{\psi_{R}}$.
In the era of noisy quantum computing without error correction, hybrid quantum-classical algorithms have been proposed for a range of applications, for example: efficiently finding the ground state energy of non-linear mean-field models [1, 2]; dynamical mean field theory algorithms for strongly correlated systems [3, 4]; and optimization problems from adjacent fields [5, 6, 7, 8, 9, 10, 11]. Two general approaches underlying several of these algorithms are the Variational Quantum Eigensolver (VQE) [12] and Quantum Subspace Expansion (QSE) [13, 14, 15] methods. QSE is a quantum-classical hybrid algorithm like VQE, but the classical optimisation problem is mapped to a generalised eigenvalue problem, which has various practical advantages compared with the highly non-linear optimisation required in VQE. In addition, it has been shown [16] that QSE can be combined with powerful error mitigation techniques [17, 18, 19], which makes this approach very attractive for noisy quantum devices.
The crucial task in QSE consists of generating a subspace with a lowest energy state that has properties which are sufficiently close to those of the true ground state. Most schemes for generating this subspace aim to create a Krylov subspace due to the excellent performance of this approach in classical algorithms [20]. A Krylov subspace is defined by a set of non-orthogonal basis states constructed through application of powers of the Hamiltonian to a reference state. An approximate Krylov subspace can be generated by performing real [13, 21, 22] and imaginary [23] time evolution, by forming a superposition of real time evolutions [24] (which can remove the need for the Hadamard test [25]), by decomposing Hamiltonian powers into products of unitary operators to form a fine grained Krylov subspace [26], or by generating Gaussian powers of the Hamiltonian [27]. Recently, it has been shown [28] that QSE with an exact Krylov basis can be realized via qubitization [29] and applied to problems from condensed matter physics, quantum chemistry and quantum field theory [28, 30].
An advantageous feature of QSE with a Krylov basis is that the accuracy of the ground state energy increases exponentially with the number of basis states [20, 31]. However, in practice the basis size cannot be increased arbitrarily because higher-order Krylov states become almost linearly dependent due to properties of the power iteration. Consequently, the overlap and Hamiltonian matrices become ill-conditioned such that noise in the matrix elements measured on quantum hardware leads to numerical instability. It has been recently shown that this problem can be mitigated by projecting the generalised eigenvalue problem onto a subspace with dominant singular values [32]. This thresholded QSE (TQSE) approach results in numerical stability up to a critical basis length, but increasing the accuracy by extending the basis size remains a desirable yet elusive capability.
Here we show that the accuracy of QSE with statistical noise can be increased beyond that achieved by TQSE while maintaining numerical stability. We achieve this by introducing a Partitioned Quantum Subspace Expansion (PQSE), where one large QSE instance is partitioned into several smaller QSE problems which exhibit better conditioning. The introduced PQSE scheme is iterative, with the output of one small QSE instance serving as input to the next small QSE problem. The scheme is illustrated in Fig. (1) for the simple case of a subspace being partitioned into two smaller subspaces [a technical introduction to the figure is provided in Sec. (2.2)]. PQSE is intended for operation in the Krylov basis regime where numerical instability becomes an obstacle to the effectiveness of TQSE. Overall PQSE requires the same quantum resources but more classical calculations than QSE or TQSE. The classical cost is determined by the choice of partitioning. By leveraging additional quantum resources, we provide a criterion for choosing an effective partitioning based on the energy variance. We analyse two canonical examples, a spin ring with random potential and a molecular Hamiltonian. We find that the size of the Krylov basis can be substantially increased compared to standard QSE and TQSE, and thus PQSE achieves substantially higher accuracy.
This paper is organised as follows. In Sec. 2.1 we introduce how a Krylov subspace is used with QSE and how thresholding can be used to recover a level of numerical stability. In Sec. 2.2 we describe the iterative subspace construction that characterises PQSE. This is followed in Sec. 3 by numerical results. In Sec. 4 we conclude with a discussion of the practical application of PQSE and further points of research, as well as a comparison with an existing iterative QSE algorithm.
2 Methods
We begin with a brief review of standard QSE and its thresholded extension TQSE in Sec. 2.1. The partitioned approach PQSE is introduced in Sec. 2.2.
2.1 Background
QSE involves diagonalising an $n$ qubit Hamiltonian $H$ in a $d$ dimensional subspace spanned by typically non-orthogonal basis states. Writing the basis states as {|$\{\ket{\phi_{i}}\}$ for $i\in\{0,\ldots d-1\}$ we can define a trial state
|$\ket{\psi}$ $=\sum\limits_{i=0}^{d-1}c_{i}\ket{\phi_{i}}\,.$ (1)
The coefficients which define the subspace ground state can be found by minimising the Ritz coefficient
$E=\frac{\braket{\psi\lvert H\rvert\psi}}{\braket{\psi}{\psi}}.$ (2)
The optimal coefficients are then solutions to the generalised eigenvalue problem
$\mathbf{H}\mathbf{c}=E\mathbf{S}\mathbf{c}\,,$ (3)
with matrix elements
$H_{ij}$ $=\braket{\phi_{i}\lvert H\rvert\phi_{j}}\,,$ (4)
$S_{ij}$ $=\braket{\phi_{i}\lvert\phi_{j}}.$ (5)
QSE will output a set of optimal coefficients $\mathbf{c}\leavevmode\nobreak\ =\leavevmode\nobreak\ [c_{0},c_{1},\dots c_{d-1}]^{T}$ which characterise the lowest energy combination of basis states with respect to $H$. The solution to Eq. (3) is a solution to the full eigenvalue problem of the Hamiltonian if and only if the variance of the Hamiltonian in this state vanishes, i.e. $\text{var}(H,\ket{\psi})=0$ with
[…]
This criterion determines a ground state solution if such a state can be expressed by the subspace basis states.
A special choice of subspace in Eq. (1) is given by a Krylov subspace. A Krylov subspace of order $R$ starting from a reference state |$\ket{\phi_{0}}$ is defined as
$\mathcal{K}_{R}(H,\ket{\phi_{0}})=\text{span}\{\ket{\phi_{0}},H\ket{\phi_{0}},\ldots,H^{R-1}\ket{\phi_{0}}\}.$ (7)
The QSE matrices are then
$H_{ij}$ $=\braket{\phi_{0}\lvert H^{i+j+1}\rvert\phi_{0}}\,,$ (8a)
$S_{ij}$ $=\braket{\phi_{0}\lvert H^{i+j}\rvert\phi_{0}}\,,$ (8b)
$i,j$$\in\{0,1,\ldots,R-1\}.$ (8c)
It has been previously noted that the QSE matrices with a Krylov basis have a Hermitian-Toeplitz structure [32]. This structure allows both QSE matrices to be constructed by calculating the first row and final column of $\mathbf{H}$.
TQSE involves projecting the QSE matrices onto the eigenspace of $\mathbf{S}$ with singular values greater than a threshold $\tau$ [32]. Concretely, if $V$ is a matrix with the eigenvectors of $\mathbf{S}$ as columns and $D$ is a diagonal matrix formed by the eigenvalues of $\mathbf{S}$ then we use $I=\{i:D_{ii}>\tau\}$ to redefine $V=V(:,I)$. $V$ projects onto the subspace containing eigenvalues above $\tau$. We solve $V^{\dagger}\mathbf{H}V\mathbf{c}=EV^{\dagger}\mathbf{S}V\mathbf{c}$ to find the lowest energy of the thresholded subspace.
2.2 Partitioned Quantum Subspace Expansion
We illustrate the idea of our algorithm by considering $P$ sequential QSE problems. A guiding schematic for $P=2$ is shown in Fig. 1. Starting from an initial state |$\ket{\phi_{0}}$, we build the subspace
$\mathcal{K}_{r_{1}}(H,\ket{\phi_{0}})\,,$ (9)
and denote its ground state by |$\ket{\psi_{1}}$. Next we form the subspace
$\mathcal{K}_{r_{2}}(H,\ket{\psi_{1}})\,,$ (10)
and denote its corresponding ground state by |$\ket{\psi_{2}}$. More generally, we consider the sequence
=$\mathcal{R}=\{r_{1},\ldots,r_{P}\},$ (11)
where each positive integer $r_{k}$ is associated with a subspace $\mathcal{K}_{r_{k}}(H,\ket{\psi_{k-1}})$ and |$\ket{\psi_{k}}$ denotes the ground state in $\mathcal{K}_{r_{k}}(H,\ket{\psi_{k-1}})$. The iterative construction implies that for all $r_{k}\in\mathcal{R}$, we have
$\mathcal{K}_{r_{k}}(H,\ket{\psi_{k-1}})\subseteq\mathcal{K}_{O(\mathcal{R})}(H,\ket{\phi_{0}})\,,$ (12)
where we define the order of the partition $\mathcal{R}$ as  [^1] It will become clear in Sec. (2.2.1) that $O(\mathcal{R})$ need not equal $R$.
$O(\mathcal{R})=1-P+\sum_{k=1}^{P}r_{k}\,.$ (13)
Consequently, each QSE problem in $\mathcal{K}_{r_{k}}(H,\ket{\psi_{k-1}})$ can be solved with the matrix elements in Eq. (2.1) required for building the subspace $\mathcal{K}_{R}(H,\ket{\phi_{0}})$. However, the ground state in $\mathcal{K}_{R}(H,\ket{\phi_{0}})$ will in general be different from the final ground state |$\ket{\psi_{P}}$ obtained by PQSE. Similarly, the ground state energies of the two approaches will be different as well. Without noise and with infinite numerical precision, the direct solution obtained by solving the QSE problem in $\mathcal{K}_{R}(H,\ket{\phi_{0}})$ is expected to be better. However, in the presence of noise and finite numerical precision this need not be the case. The advantage of PQSE is that one has to solve a sequence of smaller, better conditioned, QSE problems.
The idea of our partitioned QSE algorithm is to devise a systematic way of finding $P$ and a sequence $\mathcal{R}$ with $O(\mathcal{R})\leq R$ such that the ground state energy estimate $E_{P}$ associated with |$\ket{\psi_{P}}$ is a better approximation to the numerical estimate of the ground state energy in $\mathcal{K}_{R}(H,\ket{\phi_{0}})$. As explained in the next section, we propose to achieve this by using the variance in Eq. (6) to gauge the quality of the solution of each small QSE problem.
2.2.1 PQSE Algorithm definition
This algorithm returns an estimate $E_{g}$ of the ground state energy of a Hamiltonian $H$, as well as a description of the approximate ground state. For a given Krylov subspace order $R$, it finds a partition into smaller QSE problems $\mathcal{R}$ [see Eq. (11)] such that all small Krylov subspaces are part of $\mathcal{K}_{R}(H,\ket{\phi_{0}})$. The PQSE algorithm is as follows.
  1. 1.
    Input: Hamiltonian $H$, Krylov subspace of order $R$ with initial reference state |$\ket{\phi_{0}}$. The matrix elements defined in Eq. (2.1) are obtained using a quantum computer.
  2. 2.
    Set $r_{\text{max}}=R$, $P=0$, |$\ket{\psi_{0}}=\ket{\phi_{0}}$, $E_{g}=\bra{\psi_{0}}H\ket{\psi_{0}}$.
  3. 3.
    While ($1-P+\sum_{k=1}^{P}r_{k}<R)$ do:
    1. (a)
      Replace $P$ with $P+1$.
    2. (b)
      For all integers $q$ with $1\leq q\leq r_{\text{max}}$ do:
      1. i.
        Create the generalised eigenvalue problem in $\mathcal{K}_{q}(H,\ket{\psi_{P-1}})$ [see Appendix B].
      2. ii.
        Solve the generalised eigenvalue problem in $\mathcal{K}_{q}(H,\ket{\psi_{P-1}})$, giving coefficients $\mathbf{c}_{q}[\ket{\psi_{P-1}}]$ of the ground state estimate |$\ket{q}$ relative to the Krylov basis starting from |$\ket{\psi_{P-1}}$.
      3. iii.
        Calculate the variance $\text{var}(H,\ket{q})$ [see Appendix B.2].
    3. (c)
      Set $r_{P}=\text{argmin}_{1\leq q\leq r_{\text{max}}}\text{var}(H,\ket{q})$.
    4. (d)
      Set |$\ket{\psi_{P}}=\ket{r_{P}}$ as the ground state of the QSE problem in $\mathcal{K}_{r_{P}}(H,\ket{\psi_{P-1}})$.
    5. (e)
      if [$\text{var}(H,\ket{\psi_{P-1}})<\text{var}(H,\ket{\psi_{P}})$]:
                break
      else update
      1. $\bullet$
        $E_{g}=\bra{\psi_{P}}H\ket{\psi_{P}}/\braket{\psi_{P}}{\psi_{P}}$.
      2. $\bullet$
        $r_{\text{max}}$ with $r_{\text{max}}-(r_{P}-1)$.
  4. 4.
    Return $E_{g}$, |$\ket{\psi_{P}}$
We now quantify the resources used in PQSE. PQSE uses all matrix elements of the full Krylov subspace of order $R$. In addition, calculating the variance may require the additional matrix element $\bra{\phi_{0}}H^{2R}\ket{\phi_{0}}$, which is not needed in standard QSE, which requires the elements specified in Eq. (2.1). All QSE techniques incur the classical cost of constructing and solving Eq. (3). The additional classical cost of PQSE comes from combining the output of previous QSE problems with quantum expectation values to produce new QSE problems. While this cost may vary based on noise levels and problem parameters, we find that the classical cost scales as $\mathcal{O}(R^{4})$, see Appendix B. QSE and TQSE incur an $\mathcal{O}(R^{3})$ classical cost to solve the generalised eigenvalue problem of Eq. (3) [33].
2.3 Noise model
We now describe how we model noise on each matrix element in Eq. (2.1). Noisy versions of $\mathbf{H}$ and $\mathbf{S}$ are constructed as
$\tilde{\mathbf{H}}$ $=\mathbf{H}+\mathbf{\Delta}_{H}\,,$ (14a)
$\tilde{\mathbf{S}}$ $=\mathbf{S}+\mathbf{\Delta}_{S}\,,$ (14b)
where $\mathbf{\Delta}_{H}$ and $\mathbf{\Delta}_{S}$ are Hermitian-Toeplitz matrices with elements $\Delta_{H,ij}$ and $\Delta_{S,ij}$ which quantify the uncertainty in $\bra{\phi_{0}}H^{i+j+1}\ket{\phi_{0}}$ and $\bra{\phi_{0}}H^{i+j}\ket{\phi_{0}}$. Because $\mathbf{H}$ contains all elements of $\mathbf{S}$ except $S_{00}=\braket{\phi_{0}\lvert\phi_{0}}=1$, we only consider constructing $\mathbf{\Delta}_{H}$. We draw $\mathbf{\Delta}_{H,ij}$ from a normal distribution of width $\sigma_{H,ij}$ where
$\sigma_{H,ij}=\delta\sqrt{\bra{\phi_{0}}H^{2(i+j+1)}\ket{\phi_{0}}-\bra{\phi_{0}}H^{i+j+1}\ket{\phi_{0}}^{2}}\,,$ (15)
for a dimensionless parameter $\delta$ which controls noise strength. By considering measuring each matrix element $1/\delta^{2}$ times we see that this choice of noise model is equivalent to assuming we have access to unbiased estimators of $H_{ij}$, with any error resulting from finite sampling or limited numerical precision. This assumption can be motivated by considering applying this technique in either i) the regime where quantum error mitigation can still be effectively applied, but classical simulation is not possible [34]; ii) the regime of fault tolerant quantum computing. In [28] it is suggested that Krylov basis QSE may be a favourable choice over Phase Estimation in the early fault-tolerant regime, highlighting the former’s noise robustness as motivation.
When implementing TQSE to mitigate the effects of noise, we follow [32] and set our thresholding parameter $\tau$ to $\mathcal{O}\left((\eta^{2}_{H}+\eta^{2}_{S})^{1/2}\right)$ where $\eta_{H},\eta_{S}$ are the spectral norms of the perturbations $\mathbf{\Delta}_{H},\mathbf{\Delta}_{S}$.
3 Results
We now provide two example problems to enable a comparison between the different QSE techniques we have presented. The first example is the $n$ qubit periodic Heisenberg model with a random magnetic field in the $z$ direction, see Fig. 2(a) for a schematic and Fig. 2(b, c) for results. This is modelled by the Hamiltonian
$H_{\text{ring}}=\sum_{i}^{n}J\boldsymbol{\sigma}_{i}\cdot\boldsymbol{\sigma}_{i+1}+h_{i}\sigma^{z}_{i},$ (16)
[…]
$H_{\text{H}_{6}}=\sum_{i,j}h_{ij}a_{i}^{\dagger}a_{j}+\frac{1}{2}\sum_{i,j,k,l}h_{ijkl}a_{i}^{\dagger}a_{j}^{\dagger}a_{k}a_{l}\,.$ (17)
The fermionic creation and annihilation operators $a_{i}^{\dagger}$ and $a_{i}$ act on the $i$th spin orbital and obey $\{a_{i},a_{j}\}=0$, $\{a_{i}^{\dagger},a_{j}^{\dagger}\}=0$ and $\{a_{i},a_{j}^{\dagger}\}=\delta_{ij}\mathbf{I}$ where $\{A,B\}=AB+BA$. $h_{ij}$ and $h_{ijkl}$ are one and two electron overlap integrals respectively. We construct the molecule in the STO-3G basis [38], in equilibrium configuration at a bond length of 1.5  Å, and use a parity mapping [39] to convert from a fermionic to a 10 qubit Hamiltonian. We construct |$\ket{\phi_{0}}$ as the Hartree-Fock state. Molecular properties were computed using the PySCF library [40]. The strong electron correlation which arises in this system can prove challenging for classical methods to simulate as additional Hydrogen atoms are considered [13].
[…]
We begin with a comparison of the different QSE techniques in the absence of noise ($\delta=0$). For each Hamiltonian we vary the Krylov basis order $R$ and compute the relative error in the ground state energy
$\epsilon_{\text{rel}}=\left|\frac{E_{g}-\tilde{E}_{g}}{\tilde{E}_{g}}\right|\,,$ (18)
where $\tilde{E}_{g}$ is the ground state energy obtained by exact diagonalization and $E_{g}$ is obtained by either QSE, TQSE or PQSE.
The results for $\epsilon_{\text{rel}}$ as a function of $R$ in case of the spin ring are shown in Fig. 2(b). We find that the relative error initially decreases exponentially with $R$ for all three techniques. In spite of the absence of intentionally added noise, finite machine precision introduces a barrier to arbitrarily low solution error and results in numerical instability. In order to mitigate the effect of noise due to finite machine precision we fix the TQSE threshold parameter to $\tau=10^{-13}$, a value found to be effective in other work [28]. Note that thresholding is not used as part of PQSE.
With this, TQSE allows one to explore larger values of $R$ and the relative error is approximately one order of magnitude smaller compared to QSE. PQSE can have a slighter higher average error at basis orders where QSE fails and TQSE is numerically stable. On the other hand, PQSE returns more accurate solutions in regions where the other techniques fail. We find that by increasing $R$, PQSE can lower $\epsilon_{\text{rel}}$ by two orders of magnitude compared to TQSE.
The results for $\epsilon_{\text{rel}}$ as a function of $R$ in case of the $\text{H}_{6}$ molecule are shown in Fig. 2(e). We observe that $\epsilon_{\text{rel}}$ decays much slower with $R$ as compared to the spin ring for all QSE techniques. This finding is consistent with the results in [28], where the slow convergence with $R$ was explained by the use of a Hartree-Fock state as a reference state. QSE and TQSE display similar qualitative behaviour to the spin ring example. In contrast to the spin ring example, PQSE does not display signs of numerical instability, instead converging to a fixed value of $\epsilon_{\text{rel}}$. This fixed value is higher than the minimum error achieved by QSE and TQSE. We obtain insight into this by computing the overlap of each QSE solution with the true ground state found using exact diagonalisation. At $R=6$, QSE returns a lower energy error than PQSE, but the QSE ground state estimate has a smaller overlap with the true ground state than the PQSE estimate due to numerical instabilities. On the other hand, the overlap of the ground state estimate obtained by TQSE for $R=6$ and $R=7$ is larger than the corresponding values for PQSE, which is consistent with the better energy estimates of TQSE at these points.
Next we investigate the performance of the different QSE techniques at a fixed noise strength $\delta=10^{-6}$. Our performance metric is denoted by $\overline{\epsilon_{\text{rel}}}$, where the overbar denotes the average of $\epsilon_{\text{rel}}$ in Eq. (18) over noise instances. The corresponding standard error for $\overline{\epsilon_{\text{rel}}}$ is indicated by error bars in Figs. 2(c) and (f). In contrast to the $\delta=0$ example, we see that in both examples numerical instability restricts use of QSE to Krylov basis orders of 1 or 2. Implementing thresholding allows an exponentially fast decrease in average error for a limited range of $R$. If $R$ is increased beyond values where thresholding also exhibits instability, PQSE grants useful access to larger Krylov basis sizes and more accurate energy estimates. For the spin ring and molecular examples, in Figs. 2(c) and (f) respectively, this corresponds to a two order and half an order of magnitude decrease in $\overline{\epsilon_{\text{rel}}}$. As in the noiseless case, we find that QSE methods are more effective for the spin ring example than for $\text{H}_{6}$. Additional results for a range of noise strengths are provided in Appendix A and show similar qualitative behaviour of the relative energy error as for $\delta=10^{-6}$.
In order to investigate the performance of each approach to QSE as a function of noise strength, we compute $\overline{\epsilon_{\text{rel}}}$ for a fixed noise strength and find its minimum value as a function of $R$,
$\xi=\min\limits_{R}\overline{\epsilon_{\text{rel}}}(R)\,.$ (19)
Without noise and infinite numerical precision this minimum would be lower bounded by $\xi=0$, as one could increasing the subspace dimension indefinitely, subject to the available quantum resources. However, with noise and finite numerical precision, there will be a limit to the size of Krylov subspace dimension that can be reliably used. Therefore, this minimum provides an indicator of the lowest possible average error achieved by the considered QSE techniques.
The results for $\xi$ as a function of noise strength are shown in Figs. 3(a) and (b) for the spin ring and the $\text{H}_{6}$ molecule, respectively. We find that the the ability of PQSE to achieve a lower error on average is reflected across the range of noise strengths we consider. The reduction in error provided by PQSE is the greatest in the spin ring example, where $\xi$ can approximately two orders of magnitude below both QSE and TQSE, see Fig. 3(a). The results for $\text{H}_{6}$ in Figure 3(b) show that PQSE also provides an advantage for varying noise strengths, achieving up to an order of magnitude increase in precision.
In order to both better understand the spin ring example and the behaviour of PQSE with different choices of reference state, we apply each QSE algorithm to solving Eq. (16) for a range of values of $J/h$. We use the $J=0$ solution as an initial state |$\ket{\phi_{0}}$ for $J/h=0.1,0.2$ and the $h=0$ solution for larger ratios. The results for $\xi$ are shown in Fig. 3(c) where, for clarity, we include two noise strengths $\delta=10^{-6},10^{-1}$. We see that PQSE provides a lower minimum average error than QSE and TQSE for nearly all values of $\delta$ and all $J/h$. For $\delta=10^{-1}$ QSE has substantially larger $\xi$ than other techniques for all $J/h$. The difference between PQSE and TQSE ranges from a half order to just over one order of magnitude. For $\delta=10^{-6}$ the difference in $\xi$ between PQSE and TQSE ranges from around two orders of magnitude at $J/h=0.1$, reducing from one and a half orders to just under a half order from $J/h=0.2\dots 2$ before reaching zero difference at $J/h=5$. At $J/h=10$ TQSE outperforms PQSE by half an order of magnitude. All techniques display greater variation in $\xi$ across the range of $J/h$ for weaker noise compared with the values of $\xi$ for stronger noise. This suggests that for weaker noise the numerical stability of each QSE technique, particularly PQSE, is more sensitive to the properties of $H$ and the choice of initial state |$\ket{\phi_{0}}$, than for stronger noise.
[…]
As noted in Sec. 2.2.1, the order $O(\mathcal{R})$ of the partition $\mathcal{R}$ found by PQSE may be smaller than the maximally allowed Krylov basis order $R$. We now quantify how frequently $O(\mathcal{R})$ is smaller than $R$ and analyse the pattern of values of $O(\mathcal{R})$. Figure 4 shows the distribution of $O(\mathcal{R})$ for increasing $R$ for the spin ring example at fixed noise strength. The figure illustrates the relationship between $O(\mathcal{R})$ and $R$ as the latter increases. We see that the proportion of instances displaying early termination increases as $R$ increases and that the range of basis orders at early termination also increases. The smallest maximum allowed target Krylov basis order in this figure $R=7$, was chosen as the order at which TQSE begins to display numerical instability on average, identifiable at the minimum $\overline{\epsilon_{\text{rel}}}$ for TQSE in in Fig. 2(c). As the target basis order for PQSE increases, a large proportion of instances terminate with a basis order above this minimum. Figures 2 and 3 show that PQSE can continue increasing accuracy at basis orders above this minimum. This suggests that this implementation of PQSE is capable of detecting and avoiding numerical instabilities by choosing a stable basis sequence or terminating early.
Figure 4: Distribution of the Krylov basis order of the final basis returned by PQSE $O(\mathcal{R})$ for increasing maximum allowed Krylov basis order $R$. Data for each distribution is obtained over 100 instances of a noisy 10 qubit periodic disordered Heisenberg model. Noise is drawn from Gaussian distributions width proportional to $\delta=10^{-6}$. The maximum Krylov basis order is the far right bin, coloured in red.
4 Discussion
We have presented a classical post-processing technique that, on average and for the systems under consideration, allows the QSE algorithm to be effectively utilised with a higher dimension Krylov subspace than otherwise possible. The practical implication of this is that it can enable more accurate ground state energy estimates at the cost of additional classical processing. These improvements can be substantial: enabling two orders of magnitude improvements in accuracy. Although PQSE could get closer to the true ground state energy than the other techniques, it converged slightly slower with the Krylov basis size than the other techniques considered.
The two presented examples probed the accuracy of each QSE variant in the presence of finite sampling error. In particular, we applied all three QSE approaches to estimating the ground state energy of a quasi-random spin system and a molecular Hamiltonian. We found that in both examples, PQSE obtained more accurate energy estimates on average for the majority of levels of intentionally added noise. This advantage was more pronounced for the spin system at smaller ratios of $J/h$ than for larger ratios or the molecular system. The range of minimum average errors achieved by all techniques in the spin ring example with $J/h=0.1$ and the molecular system were substantially different. This highlights that, while PQSE is able to construct sequences of eigenvalue problems with better conditioning, and thus to outperform TQSE and QSE, it is unable to overcome limitations in accuracy set by a poor choice of initial state. Improvements to the initial state, or techniques to augment the standard Krylov basis [see Eq. (7)] with improved convergence will therefore be highly valuable. The iterative construction we have proposed leaves considerable freedom to implement these improvements. For example, one could vary the criteria used to evaluate good sequences throughout the algorithm, perhaps by initially rewarding exploration before transitioning to a more targeted criterion such as the variance. Alternatively, one could consider constructing composite subspace sequences, employing the Krylov subspace when most effective.
In order to place our new approach in context we briefly discuss PQSE in relation to an existing iterative approach to QSE, the Iterative Quantum Assisted Eigensolver (IQAE) [26]. IQAE proceeds by using the $M$ term unitary decomposition, $H=\sum_{i=1}^{M}c_{i}U_{i}$, to decompose powers of the target Hamiltonian $H$ as
$H^{k}\ket{\phi_{0}}$ $=\left(\sum_{i=1}^{M}c_{i}U_{i}\right)^{k}\ket{\phi_{0}}$ (20)
=$=\sum_{\{i_{j}\}_{j=1}^{k}=1}^{M}c_{i_{1}}\dots c_{i_{k}}U_{i_{1}}\dots U_{i_{k}}\ket{\phi_{0}}.$ (21)
This decomposition is used to form a basis as $\mathcal{B}_{k}=\{\ket{\phi_{0}}\}\cup\{U_{i}\ket{\phi_{0}}\}_{i=1}^{M}\cup\dots\cup\{U_{i_{1}}U_{i_{2}}\dots U_{i_{k}}\ket{\phi_{0}}\}_{i_{1},i_{2},\dots,i_{k}=1}^{M}$, to be contrasted with Eq. (7). Using $\mathcal{B}_{k}$ as a basis for QSE avoids the need to directly embed $H$ during a quantum computation. The lowest energy weighting of elements in $\mathcal{B}_{k}$ is computed at increasing $k\geq 0$ until a desired convergence condition is met.
Both IQAE and PQSE are iterative subspace methods, but the respective subspaces are constructed differently. Our approach is motivated by the work in [28] which presents an efficient algorithm to construct an effectively exact Krylov subspace. This is to be contrasted with the unitary decomposition used by IQAE. Furthermore, whereas IQAE works with a fixed reference state at each iteration, our procedure uses the solution from a previous iteration as a new effective reference state for the next.
A special feature of IQAE is that the basis $\mathcal{B}_{k}$ is constructed from unitaries which do not conserve the symmetry of $H$. Therefore, if the initial state |$\ket{\phi_{0}}$ lies in a different symmetry sector to the ground state of $H$, then optimising using $\mathcal{B}_{k}$ can lower the energy of |$\ket{\phi_{0}}$ while a Krylov basis constructed with a symmetry-preserving Hamiltonian $H$ would fail. An interesting avenue for future research is to consider applying PQSE with the basis used with the IQAE algorithm, to see if the dual benefits of numerical stability and symmetry breaking can be combined. Further improvements of PQSE could include augmenting the variance criterion for choosing a good sequence of partitions, and by investigating how PQSE may be more tightly integrated with thresholding.
A strategy for applying PQSE to a system where exact diagonalisation is not possible would be to gradually increase the Krylov basis size until convergence is reached, or the markers of numerical instability appear. In particular, it is beneficial to apply PQSE together with standard QSE and TQSE methods in order to increase confidence in the obtained results. Since PQSE can be implemented using the same matrix elements required for QSE and TQSE, with an additional element required if the proposed variance minimisation criterion is used, this can be done with a small increase in quantum resources. Insight into the ability of PQSE to operate at classically intractable problem sizes can be found in a recent study of the application of QSE to solving Lattice Gauge Theories [30].
Recent experimental demonstrations of quantum error correction [41, 42, 43, 44] provide some of the strongest evidence so far that the technological obstacles to fault tolerant quantum computation can be overcome. However, even the accuracy of early fault tolerant quantum computers may be limited by the impact of finite numerical precision and statistical noise on the output of algorithms such as QSE. Our approach suggests that this can be substantially alleviated.
5 Code Availability
Code for our implementation of PQSE can be found at https://github.com/t1491625/pqse_demo, where Fig. 2(c) is reproduced as a guiding example.
6 Acknowledgements
TO acknowledges support by the EPSRC through an EPSRC iCASE studentship award in collaboration with IBM Research and the Laboratory Directed Research and Development (LDRD) program of Los Alamos National Laboratory (LANL) under project number 20230049DR. DJ acknowledges support by the European Union’s Horizon Programme (HORIZON-CL4-2021-DIGITALEMERGING-02-10) Grant Agreement 101080085 QCFD. The authors would like to acknowledge the use of the University of Oxford Advanced Research Computing (ARC) facility in carrying out this work [45].
Appendix A Additional numerical results
Here we provide additional data illustrating the average relative error with increasing maximum allowed Krylov basis order associated with different approaches to QSE. We present a wider range of noise strengths $\delta$ than considered in the main text. Additional results for the spin ring example are in Fig. 5 and Hydrogen molecule example in Fig. 6. A detailed description of the setup for these examples can be found in Sec. (3) in the main text. We highlight that the results shown in Fig. 3 in the main text can be recognised by taking the minimum average error across all Krylov basis orders for each noise strength $\delta>0$. Fig. 3a connects to Fig. 5 and Fig. 3b to Fig. 6.
Figure 5: Average relative error in estimate of ground state energy for QSE with a Krylov basis: unmodified (QSE), thresholded (TQSE), partitioned (PQSE) with increasing $\delta$. The target system is the 10 qubit periodic disordered Heisenberg model with $J=0.1$ and $h=1$ for increasing Gaussian noise strength.
[…]
Appendix B Updating the reference state
Our algorithm uses the lowest energy state of one PQSE iteration as a reference state for the next. Here we show how the need to physically prepare a new reference state at each iteration can be circumvented through classical post-processing. We will show how PQSE can be implemented with only the ability to prepare the reference state |$\ket{\phi_{0}}$ which is used to construct the order $R$ Krylov subspace
|$\ket{\phi_{k}}=H^{k}\ket{\phi_{0}},\quad k\in\{0,\ldots,R-1\}\,.$ (22)
The key observation is that the matrix elements required to set up the generalised eigenvalue problem at the $P$th iteration can be expanded as a sum of the form
$H^{\{P\}}_{ij}$ $=\sum_{m=1}^{N_{H}}o_{m}\bra{\phi_{0}}H^{i+j+m}\ket{\phi_{0}}\,,$ (23)
$S^{\{P\}}_{ij}$ $=\sum_{m=0}^{N_{S}}o_{m}\bra{\phi_{0}}H^{i+j+m}\ket{\phi_{0}}\,,$ (24)
where $o_{m}$ are updated at each iteration and $11\leq N_{H}\leq 2R-$ 1 and $20\leq N_{S}\leq 2R-$ 2. To illustrate this we consider obtaining the solution |$\ket{\psi_{1}}$ from the first iteration, and the problem of constructing a new overlap matrix $S^{\{2\}}_{ij}$ for the next (with a similar derivation required to construct $H^{\{2\}}_{ij}$)
$S^{\{2\}}_{ij}$ $=\bra{\psi_{1}}H^{i+j}\ket{\psi_{1}}\,,$
$=\sum_{k,l=0}^{r_{1}-1}c^{\text{opt},*}_{r_{1},k}c^{\text{opt}}_{r_{1},l}\bra{\phi_{0}}H^{i+j+k+l}\ket{\phi_{0}}\,,$
where $c^{\text{opt}}_{r_{1},l}$ are elements of the vector of optimal QSE coefficients $\textbf{c}^{\text{opt}}_{r_{1}}[\ket{\phi_{0}}]$. For a given $20\leq m\leq 2r_{1}-$ 2 values of $k$, $l$ satisfying $k+l=m$ will contribute multiple terms with the same expectation $\bra{\phi_{0}}H^{i+j+m}\ket{\phi_{0}}$. We can group these terms
$S^{\{2\}}_{ij}$ $=\sum_{m=0}^{2r_{1}-2}\left(\sum_{\begin{subarray}{c}k,l=0:\\ k+l=m\end{subarray}}^{r_{1}-1}c^{\text{opt}*}_{r_{1},k}c^{\text{opt}}_{r_{1},l}\right)\bra{\phi_{0}}H^{i+j+m}\ket{\phi_{0}}\,,$ (27)
$=\sum_{m=0}^{2r_{1}-2}o_{m,r_{1}}\bra{\phi_{0}}H^{i+j+m}\ket{\phi_{0}}\,,$ (28)
where
$o_{m,r_{1}}=\left(\sum_{\begin{subarray}{c}k,l=0:\\ k+l=m\end{subarray}}^{r_{1}-1}c^{\text{opt}*}_{r_{1},k}c^{\text{opt}}_{r_{1},l}\right).$ (31)
Having shown how the inner product coefficients appear in Eq. (23), we now show how to update these coefficients after another iteration. If the QSE problem defined by $S^{\{2\}}_{ij}$ gives a solution |$\ket{\psi_{2}}=\sum_{k=0}^{r_{2}-1}c^{\text{opt}}_{k}H^{k}\ket{\psi_{1}}$, then at the next iteration
$S^{\{3\}}_{ij}$ $=\bra{\psi_{2}}H^{i+j}\ket{\psi_{2}}\,,$ (32)
$=\sum_{k,l=0}^{r_{2}-1}c^{\text{opt}*}_{r_{2},k}c^{\text{opt}}_{r_{2},l}\bra{\psi_{1}}H^{i+j+k+l}\ket{\psi_{1}}\,,$ (33)
$=\sum_{y=0}^{2r_{2}-2}\left(\sum_{\begin{subarray}{c}k,l=0:\\ k+l=y\end{subarray}}^{r_{2}-1}c^{\text{opt}*}_{r_{2},k}c^{\text{opt}}_{r_{2},l}\right)\sum_{m=0}^{2r_{1}-2}o_{m,r_{1}}\bra{\phi_{0}}H^{i+j+y+m}\ket{\phi_{0}}\,,$ (36)
$=\sum_{y=0}^{2r_{2}-2}\sum_{m=0}^{2r_{1}-2}o_{y,r_{2}}o_{m,r_{1}}\bra{\phi_{0}}H^{i+j+y+m}\ket{\phi_{0}}\,,$ (37)
where
$o_{y,r_{2}}=\left(\sum_{\begin{subarray}{c}k,l=0:\\ k+l=y\end{subarray}}^{r_{2}-1}c^{\text{opt}*}_{r_{2},k}c^{\text{opt}}_{r_{2},l}\right).$ (40)
Again, for a given $40\leq m^{\prime}\leq 2r_{1}+2r_{2}-$ 4 values of $y,m$ satisfying $y+m=m^{\prime}$ will contribute multiple terms with the same expectation. Grouping these terms gives
$S^{\{3\}}_{ij}$ $=\sum_{y=0}^{2r_{2}-2}\sum_{m=0}^{2r_{1}-2}o_{y,r_{2}}o_{m,r_{1}}\bra{\phi_{0}}H^{i+j+y+m}\ket{\phi_{0}}\,,$ (41)
$=\sum_{m^{\prime}=0}^{2r_{2}+2r_{1}-4}\left(\mathop{\sum_{y=0}^{2r_{2}-2}\sum_{m=0}^{2r_{1}-2}}_{y+m=m^{\prime}}o_{y,r_{2}}o_{m,r_{1}}\right)\bra{\phi_{0}}H^{i+j+m^{\prime}}\ket{\phi_{0}}\,,$ (42)
$=\sum_{m^{\prime}=0}^{2r_{2}+2r_{1}-4}o_{m^{\prime},\{r_{1},r_{2}\}}\bra{\phi_{0}}H^{i+j+m^{\prime}}\ket{\phi_{0}}\,,$ (43)
where
$o_{m^{\prime},\{r_{1},r_{2}\}}$ $=\mathop{\sum_{y=0}^{2r_{2}-2}\sum_{m=0}^{2r_{1}-2}}_{y+m=m^{\prime}}o_{y,r_{2}}o_{m,r_{1}}.$ (44)
This applies to constructing $\mathbf{S}^{\{P\}}$ at an arbitrary PQSE step as Eq. (44) can be used at each iteration after the first to update inner product coefficients $\{o_{m}\}$ using the optimal coefficients output from the most recent QSE problem $\textbf{c}^{\text{opt}}_{r_{P}}[\ket{\psi_{P-1}}]$. With each iteration, more inner products are used to construct new QSE problems. At the $P$th iteration there will be $N_{P}$ non-zero inner product coefficients where
$N_{P}=2\sum_{i=1}^{P}r_{i}-2P+1\,.$ (45)
In general, at the $P+1$th iteration the inner-product coefficients are updated as $\{o_{m,\{\mathcal{R}\}}\}_{m=0}^{N_{P-1}}\rightarrow\{o_{m^{\prime}\{\mathcal{R}^{\prime}\}}\}_{m^{\prime}=0}^{N_{P}}$ where =$\mathcal{R}=\{r_{1},r_{2},\dots,r_{P-1}\}$ and =$\mathcal{R^{\prime}}=\{r_{1},r_{2},\dots,r_{P}\}$ specify a sequence of Krylov basis orders. If the optimal QSE coefficients obtained at the $P$th iteration were $\textbf{c}^{\text{opt}}_{r_{P}}[\ket{\psi_{P-1}}]$, then the inner-product coefficients are updated using
$o_{m^{\prime},\mathcal{R}^{\prime}}$ $=\mathop{\sum_{y=0}^{2r_{P}-2}\sum_{m=0}^{N_{P-1}}}_{y+m=m^{\prime}}o_{y,r_{P}}o_{m,\mathcal{R}}\,,$ (46)
where
$o_{y,r_{P}}=\left(\sum_{\begin{subarray}{c}k,l=0:\\ k+l=y\end{subarray}}^{r_{P}-1}c^{\text{opt}*}_{r_{P},k}c^{\text{opt}}_{r_{P},l}\right).$ (49)
As PQSE has the constraint that
$\sum_{i=1}^{P}r_{i}-P+1\leq R\,,$ (50)
there will be at most $N_{P}=2R-1$ inner product coefficients. If we consider $S_{ij}=\bra{\phi_{0}}H^{i+j}\ket{\phi_{0}}$ for single-step QSE with $10\leq i,j\leq R-$ 1, we see $N_{P}$ matches the number of unique inner products required to construct $\mathbf{S}$.
B.1 Resource cost
Here we estimate the classical cost of PQSE. Our starting point is the observation that any sub-QSE problem with reference state |$\ket{\psi}$ at the $P>1$th PQSE iteration requires constructing matrices $\mathbf{H}^{\{P\}}$ and $\mathbf{S}^{\{P\}}$. The latter can be written
$S^{\{P\}}_{ij}$ $=\bra{\psi}H^{i+j}\ket{\psi}\,,$ (51)
$=\sum_{m=0}^{N_{P-1}}o_{m}\bra{\phi_{0}}H^{i+j+m}\ket{\phi_{0}}\,,$ (52)
where $20\leq N_{P}\leq 2R-$ 2 and $N_{P}$ is defined in Eq. (45). The cost of constructing $S^{\{P+1\}}_{ij}$ at a new iteration comes from the cost of updating the inner-product coefficients $\{o_{m,\mathcal{R}}\}_{m=0}^{N_{P-1}}\rightarrow\{o_{m^{\prime},\mathcal{R}^{\prime}}\}_{m^{\prime}=0}^{N_{P}}$. If the subspace dimension chosen at the $P$th iteration is $r_{P}$, then the coefficients are updated using Eqs. (46) and (49). Computing $\{o_{y,r_{P}}\}_{y=0}^{2r_{P}-2}$ in Eq. (49) can be done in $\mathcal{O}(r_{P}^{2})$ basic arithmetic operations. Computing $\{o_{m^{\prime},\mathcal{R}^{\prime}}\}_{m^{\prime}=0}^{N_{P}}$ in Eq. (46) requires $\mathcal{O}(r_{P}N_{P-1})$ basic arithmetic operations. The cost of updating the inner-product coefficients is therefore at most $\mathcal{O}(R^{2})$.
With the updated coefficients in hand, constructing an element of $\mathbf{S}^{\{P\}}$ using Eq. (52) requires $N_{P}$ basic steps. The same inner-product coefficients can be used to construct a single element of $\mathbf{H}^{\{P\}}$, which also requires $N_{P}$ basic steps. By making use of the Hermitian-Toeplitz structure of the $\mathbf{S}$ and $\mathbf{H}$ matrices, setting up the generalised eigenvalue problem in $\mathcal{K}_{r_{P+1}}(H,\ket{\psi})$ requires $4r_{P+1}N_{P}$ operations. Constructing all possible subspaces for a given $r_{P+1}$ and solving the generalised eigenvalue problem with cost $wr^{3}$ [33], where $w$ is a constant factor, has thus the overall cost
$C(P)\leq R^{2}+\sum\limits_{r_{P+1}=2}^{R-A}(4r_{P+1}N_{P}+wr^{3})\,,$ (53)
Where $A=\sum_{i=1}^{P}r_{i}-P+1$ limits $r_{P+1}$ in a way that satisfies the constraint in Eq. (50). Evaluating the sum shows that the dominant term scales like $R^{4}$. A full run of PQSE returns the sequence =$\mathcal{R}=\{r_{1},\ldots,r_{B}\}$ and thus requires $B$ times the steps detailed above. It follows that the numerical cost for the classical part of PQSE is upper-bounded by $BR^{4}$.
B.2 Calculating expectation values
PQSE returns |$\ket{\psi_{P}}$ as an approximation to the ground state of $H$ in terms of expansion coefficients $\textbf{c}^{\text{opt}}_{r_{P-1}}[\ket{\phi_{P-1}}]$ relative to the Krylov basis with starting vector |$\ket{\psi_{P-1}}$. We obtain the energy of this state $E_{P}$ when solving Eq. (3) in $\mathcal{K}_{r_{P}}(H,\ket{\psi_{P-1}})$ [see Sec. 2.1 and 2.2 in the main text]. In this section we show how to compute expectation values of observables in the state |$\ket{\psi_{P}}$ in terms of expectation values in the initial state |$\ket{\phi_{0}}$. For example, this allows us to calculate the energy variance in |$\ket{\psi_{P}}$.
The expected value of observable $\hat{O}$ after one PQSE iteration can be written
$\bra{\psi_{1}}\hat{O}\ket{\psi_{1}}$ $=\sum_{k,l=0}^{r_{1}-1}c^{\text{opt},*}_{r_{1},k}c^{\text{opt}}_{r_{1},l}\bra{\phi_{0}}H^{k}\hat{O}H^{l}\ket{\phi_{0}}\,,$ (54)
and after another PQSE iteration
$\bra{\psi_{2}}\hat{O}\ket{\psi_{2}}$ $=\sum_{k_{2},l_{2}=0}^{r_{2}-1}c^{\text{opt},*}_{r_{2},k_{2}}c^{\text{opt}}_{r_{2},l_{2}}\bra{\psi_{1}}H^{k_{2}}\hat{O}H^{l_{2}}\ket{\psi_{1}}\,,$ (55)
$=\sum_{k_{2},l_{2}=0}^{r_{2}-1}\sum_{k_{1},l_{1}=0}^{r_{1}-1}c^{\text{opt},*}_{r_{2},k_{2}}c^{\text{opt},*}_{r_{1},k_{1}}c^{\text{opt}}_{r_{2},l_{2}}c^{\text{opt}}_{r_{1},l_{1}}\bra{\phi_{0}}H^{k_{1}+k_{2}}\hat{O}H^{l_{1}+l_{2}}\ket{\phi_{0}}\,.$ (56)
For a given $20\leq m,m^{\prime}\leq r_{1}+r_{2}-$ 2 values of $k_{1}$, $k_{2}$ satisfying $k_{1}+k_{2}=m$ and values of $l_{1}$, $l_{2}$ satisfying $l_{1}+l_{2}=m^{\prime}$ will contribute multiple terms with the same expectation $\bra{\phi_{0}}H^{m}\hat{O}H^{m^{\prime}}\ket{\phi_{0}}$. We can group these terms
$\bra{\psi_{2}}\hat{O}\ket{\psi_{2}}$ $=\sum_{m,m^{\prime}=0}^{r_{1}+r_{2}-2}\left(\mathop{\sum_{k_{2}=0}^{r_{2}-1}\sum_{k_{1},=0}^{r_{1}-1}}_{k_{1}+k_{2}=m}c^{\text{opt},*}_{r_{2},k_{2}}c^{\text{opt},*}_{r_{1},k_{1}}\right)\left(\mathop{\sum_{l_{2}=0}^{r_{2}-1}\sum_{l_{1}=0}^{r_{1}-1}}_{l_{1}+l_{2}=m^{\prime}}c^{\text{opt}}_{r_{2},l_{2}}c^{\text{opt}}_{r_{1},l_{1}}\right)\bra{\phi_{0}}H^{m}\hat{O}H^{m^{\prime}}\ket{\phi_{0}}\,.$ (57)
$=\sum_{m,m^{\prime}=0}^{r_{1}+r_{2}-2}b^{*}_{m,\{r_{1},r_{2})\}}b_{m^{\prime},\{r_{1},r_{2}\}}\bra{\phi_{0}}H^{m}\hat{O}H^{m^{\prime}}\ket{\phi_{0}}\,,$ (58)
where
$b_{m^{\prime},\{r_{1},r_{2}\}}$ $=\mathop{\sum_{l_{2}=0}^{r_{2}-1}\sum_{l_{1}=0}^{r_{1}-1}}_{l_{2}+l_{1}=m^{\prime}}c^{\text{opt}}_{r_{2},l_{2}}c^{\text{opt}}_{r_{1},l_{1}}\,,$ (59)
and $b_{m^{\prime},\{r_{1},r_{2}\}}^{*}$ denotes the complex conjugate. By considering computing an observable of the output of another PQSE iteration, and grouping inner products as in Eq. (57), we learn the rule to update $b_{m^{\prime},\{r_{1},r_{2}\}}$
$\bra{\psi_{3}}\hat{O}\ket{\psi_{3}}$ $=\sum_{k_{3},l_{3}=0}^{r_{3}-1}c^{\text{opt},*}_{r_{3},k_{3}}c^{\text{opt}}_{r_{3},l_{3}}\bra{\psi_{2}}H^{k_{3}}\hat{O}H^{l_{3}}\ket{\psi_{2}}\,,$ (60)
$=\sum_{k_{3},l_{3}=0}^{r_{3}-1}c^{\text{opt},*}_{r_{3},k_{3}}c^{\text{opt}}_{r_{3},l_{3}}\sum_{k_{2},l_{2}=0}^{r_{2}-1}\sum_{k_{1},l_{1}=0}^{r_{1}-1}c^{\text{opt},*}_{r_{2},k_{2}}c^{\text{opt},*}_{r_{1},k_{1}}c^{\text{opt}}_{r_{2},l_{2}}c^{\text{opt}}_{r_{1},l_{1}}\bra{\phi_{0}}H^{k_{1}+k_{2}+k_{3}}\hat{O}H^{l_{1}+l_{2}+l_{3}}\ket{\phi_{0}}\,.$ (61)
$=\sum_{k_{3},l_{3}=0}^{r_{3}-1}c^{\text{opt},*}_{r_{3},k_{3}}c^{\text{opt}}_{r_{3},l_{3}}\sum_{m,m^{\prime}=0}^{r_{1}+r_{2}-2}b^{*}_{m,\{r_{1},r_{2}\}}b_{m^{\prime},\{r_{1},r_{2}\}}\bra{\phi_{0}}H^{m+k_{3}}\hat{O}H^{m^{\prime}+l_{3}}\ket{\phi_{0}}\,.$ (62)
$=\sum_{m^{\prime\prime},m^{\prime\prime\prime}=0}^{r_{1}+r_{2}+r_{3}-3}\left(\sum_{k_{3}=0}^{r_{3}-1}\sum_{m=0}^{r_{1}+r_{2}-2}c^{\text{opt},*}_{k_{3}}b^{*}_{m,\{r_{1},r_{2}\}}\right)\left(\sum_{l_{3}=0}^{r_{3}-1}\sum_{m^{\prime}=0}^{r_{1}+r_{2}-2}c^{\text{opt}}_{l_{3}}b_{m^{\prime},\{r_{1},r_{2}\}}\right)\bra{\phi_{0}}H^{m^{\prime\prime}}\hat{O}H^{m^{\prime\prime\prime}}\ket{\phi_{0}}\,.$ (63)
$=\sum_{m^{\prime\prime},m^{\prime\prime\prime}=0}^{r_{1}+r_{2}+r_{3}-3}b^{*}_{m^{\prime\prime},\{r_{1},r_{2},r_{3}\}}b_{m^{\prime\prime\prime},\{r_{1},r_{2},r_{3}\}}\bra{\phi_{0}}H^{m^{\prime\prime}}\hat{O}H^{m^{\prime\prime\prime}}\ket{\phi_{0}}\,,$ (64)
where
$b_{m^{\prime\prime},\{r_{1},r_{2},r_{3}\}}=\sum_{k_{3}=0}^{r_{3}-1}\sum_{m^{\prime}=0}^{r_{1}+r_{2}-2}c^{\text{opt}}_{r_{3},k_{3}}b_{m,\{r_{1},r_{2}\}}\,.$ (65)
In this way we can reconstruct properties of the PQSE output state in terms of properties of the initial Krylov basis states. The rule in Eq. (65) generalises to the $P$th PQSE iteration as
$b_{m^{\prime\prime},\mathcal{R}^{\prime}}=\sum_{k=0}^{r_{P}-1}\sum_{m=0}^{O(\mathcal{R})-1}c^{\text{opt}}_{r_{P},k}b_{m,\mathcal{R}}\,,$ (66)
where =$\mathcal{R}^{\prime}=\{r_{1},r_{2},\dots,r_{P}\}$ and =$\mathcal{R}=\{r_{1},r_{2},\dots,r_{P-1}\}$ and $O(\mathcal{R})=1-(P-1)+\sum_{k=1}^{P-1}r_{k}$ as introduced in the main text in Sec. (2.2). By iteratively constructing the coefficients $b_{m,\mathcal{R}}$ using the new optimal QSE coefficients $\textbf{c}^{\text{opt}}_{r_{P}}[\ket{\psi_{P-1}}]$ at each iteration we are able to compute the observable of the PQSE output state at the $P$th iteration as
$\bra{\psi_{P}}\hat{O}\ket{\psi_{P}}$ $=\sum_{m,m^{\prime}=0}^{O(\mathcal{R})-1}b^{*}_{m,\mathcal{R}}b_{m^{\prime},\mathcal{R}}\bra{\phi_{0}}H^{m}\hat{O}H^{m^{\prime}}\ket{\phi_{0}}\,.$ (67)